OpenStudy (justcatx3):

Express the following function, h(x) as a composition of two functions f and g. h(x) = x^2/(x^2 + 4) I've been trying to wrap my head around this, but can't seem to come to an answer. At first I wanted to say: f(x) = x^2/x g(x) = x^2 + 4 But then I realized that g(x) would have be put in both x values in f(x) which is not the function given. Any suggestions?

TheSmartOne (thesmartone):

Try with f(x) = x/(x+4) what would g(x) be? :) :P

OpenStudy (justcatx3):

I assume to get a squared answer g(x) would equal x.

TheSmartOne (thesmartone):

nope, composition of a function basically means you put one function inside of another for example if f(x) = x + 2 and g(x) = 2x + 2 then f(g(x)) = f(2x + 2) = (2x + 2) + 2 = 2x + 4 mhmm, do you understand or should I do a better example? :)

OpenStudy (justcatx3):

I understand. So then g(x) would equal x^2. So that x^2 would be inputted into the x values in f(x).

TheSmartOne (thesmartone):

Correct :)

OpenStudy (justcatx3):

Awesome! Thanks a bunch, this question has been driving me crazy, haha.

TheSmartOne (thesmartone):

Anytime! I don't know how I saw the numbers. I was just thinking and it clicked. It all comes with practice, I suppose. xD

OpenStudy (justcatx3):

I suppose so. It would have been easier for me if it was like: (2x+1)^2 xD I would have figured that would out with no problem.

TheSmartOne (thesmartone):

Anyways, welcome to OpenStudy :)

OpenStudy (justcatx3):

Thanks! OpenStudy has definitely been helpful for me so far :) Definitely helps if you are stuck.

OpenStudy (agent0smith):

You can also just cheat and go with f(x) = x g(x) = x^2/(x^2 + 4) then f(g(x)) = x^2/(x^2 + 4) but it's basically cheating, lol

TheSmartOne (thesmartone):

LOL. I can imagine the teacher's reaction if you did something like that x'd

OpenStudy (justcatx3):

Oh my gosh, I can see what that's basically cheating, haha. Right? It's like waiitt a second XD

OpenStudy (justcatx3):

Or maybe they would think your really clever, who knows? However, I prefer not to find out xD

TheSmartOne (thesmartone):

What would your reaction be, Agent, if you got an answer like that? x'd

OpenStudy (agent0smith):

I'd probably still accept it, I remember teaching this to a precalc class and thinking they still have to understand it, to do that. Less cheaty other options are ones like f(x) = (x-4)/x g(x) = x^2 + 4 f(g(x)) = x^2/(x^2+4)

TheSmartOne (thesmartone):

Nice. The possibilities are endless I'm sure. It'll just get more complicated xD

OpenStudy (agent0smith):

Actually the most cheaty option is f(x) = x^2/(x^2+4) g(x) = x f(g(x)) = f(x) = lol

OpenStudy (agent0smith):

Put on your lollerskates for that last one.

TheSmartOne (thesmartone):

g(x) = x^2/(x^2 + 4) f(x) = x g(f(x)) = g(x) it just has to be a composition of the functions, they didn't specify whether it had to be f(g(x)) or g(f(x)) ;p

OpenStudy (agent0smith):

Or go for a ride on the lollercoaster http://rianjs.net/images/lol/LOLLERCOASTER.gif

OpenStudy (justcatx3):

That gif. Is awesome xD

TheSmartOne (thesmartone):

That's the first time I've seen that lollercoaster. LOL

OpenStudy (agent0smith):

The roflcopter is an old favourite http://www.kreativekorp.com/miscpages/omgwtfbbq/roflcopter.gif

TheSmartOne (thesmartone):

Where do you find these stuff. ROFL. I've never seen this before. ROFLOL

OpenStudy (justcatx3):

Oh my gosh, haha! Brilliant.

OpenStudy (agent0smith):

Here's the lollerskates gif http://www.threadbombing.com/data/media/2/lollerskates.gif lol i googled lollerskates

OpenStudy (agent0smith):

Steamloller http://bbsimg.ngfiles.com/1/5573000/ngbbs42bd6adc5beb9.gif

TheSmartOne (thesmartone):

http://rofl.name/a/ x'd

OpenStudy (agent0smith):

x'd

TheSmartOne (thesmartone):

hey, this applies to us because of the olympics going on xDD http://rofl.name/media.php?q=aniascii_lolympics.gif x'd

OpenStudy (justcatx3):

Thanks for opening me to awesome new world x'D But I have to go. Thanks so much for the help!

TheSmartOne (thesmartone):

http://rofl.name/media.php?q=aniascii_kthxbye.gif Cya! If you ever need any help, we'll be around. Just give us a tag like @agent0smith :)

OpenStudy (justcatx3):

Ah ok! I wasn't sure how to do that. Thank you :)

TheSmartOne (thesmartone):

And, one last thing, my username is misleading. xD Agent is smarter than me in lotsa stuff. xD

OpenStudy (justcatx3):

Your icon still cool though x'D I love cats so ya xD

TheSmartOne (thesmartone):

I just recently put it up as my profile picture xD

OpenStudy (agent0smith):

@TheSmartOne your links don't work, i don't think you can link them individually? The lolrus isn't happy about not being featured on that page http://bbsimg.ngfiles.com/1/5573000/ngbbs42bd67f355412.gif

OpenStudy (agent0smith):

Also @TheSmartOne I've got like 15 years on you, and a university degree of knowledge, too ;P

TheSmartOne (thesmartone):

Just wait till I'm older than you ;)

OpenStudy (justcatx3):

Lol, cya guys!

TheSmartOne (thesmartone):

See you around (:

OpenStudy (justcatx3):

You too! Bye~

TheSmartOne (thesmartone):

doooodd x'd

OpenStudy (agent0smith):

LMFAO train http://bbsimg.ngfiles.com/1/5573000/ngbbs42bd73cb794b1.gif

OpenStudy (zzr0ck3r):

If someone sets one of the functions as the identity function, they most likely completely understand this topic. But they might have been told to just do this every time and it will 'work' so I tell my students, after making the class guess what would be the easiest way, that they are not allowed to do it.

OpenStudy (marcelie):

lmaoo x'd

OpenStudy (agent0smith):

Let's all have a sip of lmaonade while we watch these. 3D roflcopter! http://orig02.deviantart.net/7044/f/2009/240/1/6/animated_roflcopter__by_rydian.gif Non-functioning tail rotor means it couldn't fly without spinning the body spinning, due to conservation of angular momentum. Helicopters are interesting machines. A lot of smart people think they shouldn't be able to fly.